# Blockwise SFT for Diffusion Language Models: Reconciling Bidirectional Attention and Autoregressive Decoding

### v1

@@ 3.4 Practical Implementation of Blockwise SFT
1:Instruction–response pair $\mathbf{x}=[\mathbf{c};\mathbf{r}]$; block size $B$; diffusion steps $T$; per-step weights $\{\omega_{t}\}_{t=1}^{T}$; model $\theta$; forward kernel $q_{t}$
2:Partition response into blocks $\mathbf{r}=[\mathbf{b}^{(1)},\dots,\mathbf{b}^{(M)}]$ with index sets $\{\mathcal{I}_{m}\}_{m=1}^{M}$ where $M=\lceil L_{r}/B\rceil$
@@
4:Sample masking rate $\pi\sim\mathrm{Uniform}(0,1)$
5:Build mask $\mathbf{m}_{1:L}$:
  • Prefix ($i\leq L_{c}+B(a-1)$): $m_{i}\leftarrow 0$ (clean; no gradient)
  • Active block ($i\in\mathcal{I}_{a}$): $m_{i}\leftarrow\mathrm{Bernoulli}(\pi)$
  • Suffix ($i>L_{c}+Ba$): $m_{i}\leftarrow 1$ (fully hidden)
6:Sample diffusion step $t$ with probability $\tilde{\omega}_{t}=\omega_{t}/\sum_{s=1}^{T}\omega_{s}$ (single-$t$ variant; see App. A.5)
7:Noise the sequence: $\mathbf{z}_{t}\sim q_{t}(\cdot\mid\mathbf{x})$
@@
8:Compute block-local loss $\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)=-\sum_{i\in\mathcal{I}_{a}}\log p_{\theta}(x_{i}\mid\mathbf{z}_{t},t)$
9:Scale by $Z=\sum_{s=1}^{T}\omega_{s}$ and take gradient of $Z\cdot\mathbb{E}_{\mathbf{z}_{t}}[\tilde{\mathcal{L}}_{t}]$ (unbiased for equation 7)
@@
10:Backpropagate only through positions $i\in\mathcal{I}_{a}$; treat prefix as constants and suffix as masked
11:Update $\theta$ with AdamW and the shared hyperparameters (App. A.1)
4 Experiment
4.1 Datasets and Tasks
MetaMathQA.
We train on meta-math/MetaMathQA [^1] https://huggingface.co/datasets/meta-math/MetaMathQA, a large-scale math QA corpus built from GSM8K and MATH training questions via multiple rewrites, with no test leakage. The commonly used split contains $\sim$395K examples, each paired with a reference answer.
Evaluation Sets.
For evaluation, we use the test splits of openai/gsm8k [^2] https://huggingface.co/datasets/openai/gsm8k and HuggingFaceH4/MATH-500333https://huggingface.co/datasets/HuggingFaceH4/MATH-500. GSM8K contains grade-school–level arithmetic word problems, while MATH covers advanced competition-style problems across algebra, geometry, number theory, and combinatorics. We report Pass@1 accuracy, computed by exact string match between the model’s prediction and the reference answer.
4.2 Baselines and Comparisons
Models.
We compare three configurations: Base (no supervision), Classical SFT (randomly mask the full response each step), and Blockwise SFT (mask only the active block while keeping the prefix clean and the suffix hidden).
Why two comparison protocols?
The two methods differ in (i) how much compute each update costs and (ii) how many response tokens receive supervision. To make the comparison fair, we report results under two complementary controls: Equal-FLOPS (fix compute) and Equal-Tokens (fix supervised-token budget).
Equal-FLOPS.
For diffusion LMs, each training step runs the Transformer over the entire sequence at every layer (and diffusion step), so the dominant cost—self-attention and MLP—depends on sequence length, not on which tokens receive loss. Consequently, when optimizer steps, batch size, sequence length, and the number of diffusion steps are fixed, the per-step FLOPs are essentially identical for Classical SFT and Blockwise SFT; masking only changes where the loss (and nonzero gradients) appear, not the forward/backward compute graph. Example: with $L_{r}{=}128$ and B=32B{=}32 (M=4M{=}4 blocks), one epoch over $N$ samples executes the same forward/backward work for both methods; Classical SFT supervises all $128$ response tokens, whereas Blockwise SFT supervises only the $32$ tokens in the sampled block. In practice, we therefore train both for one epoch on the same $N$ with identical settings—matching compute while varying where gradients land (see Figure 2).
Equal-Tokens.
Here we match the total number of supervised response tokens. We define one traversal as: on average, each response token is supervised once. Using the same example ($L_{r}{=}128$, B=32B{=}32), after one epoch Classical SFT achieves one traversal (all $128$ tokens supervised), while Blockwise SFT achieves $1/4$ traversal (only one $32$-token block). To compare at one traversal, we run Classical SFT for 1 epoch and Blockwise SFT for 4 epochs, sampling different blocks over time. This protocol fixes the supervised-token budget and isolates how supervision is distributed across blocks; compute is not constrained.
Figure 2: Equal-FLOPS vs. Equal-Tokens.
Implementation details.
We fine-tune all models with LoRA (Hu et al., 2021). We adopt the hyperparameters of Biderman et al. (2024) verbatim, because LoRA is sensitive to optimizer and adapter settings and exposes many coupled knobs; exhaustive search is costly and unnecessary for our purpose. We scale the learning rate with the square-root rule (Krizhevsky, 2014).
For both methods, the Bernoulli masking rate is sampled as $\pi\sim\mathrm{Uniform}(10^{-3},1)$ at the sequence (Classical SFT) or active-block (Blockwise SFT) level, following the fine-tuning recipe of LLaDA (Nie et al., 2025). After applying the Bernoulli mask, if no position is masked in the supervision region, we additionally mask the last token of that region—i.e., the last token of the full response for Classical SFT, or the last token of the active block for Blockwise SFT.
Experiments use GSAI-ML/LLaDA-8B-Instruct [^4] https://huggingface.co/GSAI-ML/LLaDA-8B-Instruct. All runs use a single NVIDIA A100 (80 GB). With identical dataloader and optimizer settings, both methods reach $\sim$5.1k tokens s-1. We train in bfloat16 (Micikevicius et al., 2018) using AdamW (Loshchilov & Hutter, 2019) with linear warm-up and cosine decay (Smith, 2017). Full hyperparameters appear in Appendix A.1. With per-device batch size $1$, peak usage is $\approx\!23.5$ GB, so all experiments are feasible on a 24 GB consumer GPU (e.g., RTX 3090) or any device with comparable or larger memory.
@@ 4.3 Main Results
Results under Equal-FLOPS.
(a) Classical SFT
(b) Blockwise SFT
Figure 3: Loss as a function of training steps.
Figure 4: Performance as a function of training steps under Equal-FLOPS.
Figure 3 first presents the training loss. Both methods show a rapid drop in the first 50–100 steps, after which the exponential-moving-average (EMA) curves flatten. Near convergence the short-term oscillation amplitudes are similar, but the steady levels differ markedly: Blockwise SFT stabilizes around 0.24 (3(b)), whereas Classical SFT stabilizes around 0.36 (3(a)).
Figure 4 reports Pass@1 under matched forward–backward token operations (Equal-FLOPS) on GSM8K (4(a)) and MATH (4(b)). On GSM8K, all methods improve sharply within the first 100–200 steps and then slow, but Blockwise SFT stays well above Classical SFT throughout, reaching $\approx\!+10$ Pass@1 at the final checkpoint and consistently exceeding the LLaDA-8B-Instruct baseline. On MATH, Blockwise SFT again improves steadily, whereas Classical SFT degrades over training and remains below the Base model at every checkpoint. These trends indicate that supervision aligned with semi-autoregressive decoding yields stronger accuracy under fixed compute.
Results under Equal-Tokens.
Figure 5: Performance as a function of training steps under Equal-Tokens.
Figure 5 reports Pass@1 as a function of traversal count $\tau$ under Equal-Tokens (5(a) for GSM8K and 5(b) for MATH; cf. §4.2). Across both datasets, Blockwise SFT achieves a substantial gain already at $\tau=1$ and then remains comparatively stable as $\tau$ increases (through $\tau=3$).
On GSM8K, Classical SFT shows only a small lift at $\tau=1$, far below Blockwise SFT, and then drops rapidly; by $\tau=3$ it underperforms the Base model. On MATH, Classical SFT degrades steadily with larger $\tau$, typically below the baseline across most traversals, while Blockwise SFT makes modest gains up to $\tau=\!1$ and then stabilizes. These results mirror the Equal-FLOPS trends and indicate that block-aligned supervision delivers earlier gains and more stable learning under matched token budgets.
4.4 Head-to-Head Comparison
We benchmark seven systems under matched compute (Equal-FLOPS): the Base model, Classical SFT (D3PM-style masked diffusion (Austin et al., 2023)), four improved SFT variants—MDLM (Sahoo et al., 2024), Soft-Masked Diffusion LM (Chen et al., 2023), RDM (Zheng et al., 2024), and Two-Step Loss & Scheduling (Asada & Miwa, 2025)—and our Blockwise SFT. All models are fine-tuned on MetaMathQA with identical tokenization, optimization, and decoding settings; only the supervision mechanism differs. As summarized in Table 1, Blockwise SFT attains the best Pass@1 on both benchmarks (GSM8K: $\mathbf{76.0}\pm 1.6$; MATH: $\mathbf{34.2}\pm 0.5$). Relative to the strongest non–Blockwise baseline (Two-Step Loss & Scheduling, $70.8\pm 1.3$ / $32.6\pm 0.7$), the gains are +5.2+5.2 points on GSM8K and +1.6+1.6 on MATH. On MATH, Classical SFT trails the Base model ($29.6\pm 0.7$ vs. $31.7\pm 0.4$), while recent variants offer only modest improvements (e.g., RDM $32.3\pm 0.6$, Two-Step $32.6\pm 0.7$). On GSM8K, several variants exceed Classical SFT (e.g., MDLM $68.4\pm 1.9$), but none approach Blockwise SFT. Overall, the head-to-head results under matched compute substantiate the superiority of Blockwise SFT over 4 other enhanced algorithms derived from Classical SFT..
@@ 4.5 Block Size Consistency Study
Figure 6: Combinations of Blockwise SFT training block size $B_{\text{train}}$ and inference block size $B_{\text{infer}}$.
We train two $4\times 4$ grids over $B_{\text{train}},B_{\text{infer}}\in\{8,16,32,64\}$ to measure sensitivity to block-size mismatch. As shown in Figure 6, accuracy is highest on the diagonal ($B_{\text{train}}{=}B_{\text{infer}}$) and remains strong for near-diagonal pairs (small $|B_{\text{train}}-B_{\text{infer}}|$), while it degrades markedly as combinations move farther off the diagonal (e.g., $8{\to}64$ or $64{\to}8$). Across both heatmaps, the diagonal and its immediate neighbors form the best-performing band, indicating that Blockwise SFT’s gains are primarily driven by aligning training with the semi-autoregressive decoding granularity rather than any particular block size.
4.6 Ablation Experiments
Figure 7: Performance as a function of $\pi_{\text{prefix}}$.
Figure 8: Performance as a function of $\pi_{\text{suffix}}$.
Prefix noise
($\pi_{\text{prefix}}$, mask ratio over prefix blocks). In Figure 7, a small mask rate (e.g., $\pi_{\text{prefix}}{=}0.33$) causes only minor loss on both GSM8K and MATH, but accuracy degrades markedly as $\pi_{\text{prefix}}$ increases, with the largest drop at $\pi_{\text{prefix}}{=}1.0$. This pattern matches the mechanism: prefix masking corrupts the conditioning context except when the active block is the first (no prefix to mask), so $\pi_{\text{prefix}}{=}1$ behaves like a reduction in the fraction of instances that contribute gradients—some samples remain unaffected—but the frequent noisy contexts still hurt performance. Overall, the ablation confirms that controlling prefix contamination is a key factor for Blockwise SFT, consistent with the need for a clean, deterministic prefix at inference.
Suffix leakage
($\pi_{\text{suffix}}$, mask ratio over suffix blocks). In Figure 8, performance drops clearly as $\pi_{\text{suffix}}$ decreases (i.e., as more future tokens become visible). When $\pi_{\text{suffix}}=0$ (fully visible suffix), the scores approach those of Classical SFT on both datasets, indicating that strict future masking is the key factor behind Blockwise SFT’s advantage. Implementation details for both ablations are provided in Appendix A.7.
@@ 5 Conclusion
We introduced Blockwise SFT, a simple fine-tuning strategy that aligns supervision with semi-autoregressive, blockwise decoding in discrete diffusion LMs. By freezing clean prefixes, masking only the active block, and strictly hiding future tokens, Blockwise SFT removes the contextual shift and dependency leakage inherent to classical SFT. The objective is theoretically grounded—providing a variational upper bound on blockwise likelihoods, admitting unbiased block/time-sampled gradients, and explaining the bias of classical SFT under mismatch (Theorems 3.13.3). Empirically, fine-tuning on MetaMathQA with evaluation on GSM8K and MATH yields consistent gains under matched compute and token budgets, state-of-the-art results in a head-to-head comparison, and diagnostic trends that track training–inference alignment (block-size consistency in §4.5, prefix/suffix ablations in §4.6).
Beyond accuracy, the recipe is easy to adopt—no architectural changes, a drop-in loss, and practical single-step estimators (§3.4). Limitations and avenues for future work include adaptive or uncertainty-aware block sizing, combining blockwise supervision with preference/RL-style objectives for instruction following, and extending the alignment principle to other semi- or non-autoregressive generators. Overall, respecting the structural constraints of the decoding procedure is not just an efficiency tweak but a core ingredient for performance in diffusion-based language models.
@@ A.5 Practical Estimators and Weighting
To reduce cost, we may sample a single diffusion step $t$ with probability $\tilde{\omega}_{t}\propto\omega_{t}$ and multiply the per-step gradient by the normalizer $Z=\sum_{s=1}^{T}\omega_{s}$. This recovers an unbiased estimate of the $t$-sum in equation 4. Formally, letting
@@
which equals $\nabla_{\theta}\,\mathcal{L}_{\text{BW--SFT}}(\theta)$. Under Assumption 1 (local surrogate fidelity), replace the inner gradient by that of the t=0t{=}0 blockwise NLL to obtain $\nabla_{\theta}\,\mathcal{R}_{\text{block}}(\theta)$ up to a global scalar. ∎
Proof of Theorem 3.3.

### v2

@@ 3.4 Practical Implementation
1:Instruction–response pair $\mathbf{x}=[\mathbf{c};\mathbf{r}]$; block size $B$; diffusion steps $T$; weights $\{\omega_{t}\}_{t=1}^{T}$; model $\theta$
2:Partition response into blocks $\mathbf{r}=[\mathbf{b}^{(1)},\dots,\mathbf{b}^{(M)}]$ where $M=\lceil L_{r}/B\rceil$
@@
4:Sample mask rate $\pi\sim\mathrm{Uniform}(0,1)$
5:Construct training mask $\mathbf{m}_{1:L}$:
  • Prefix ($i\leq L_{c}+B(a-1)$): $m_{i}=0$
  • Active block ($i\in\mathcal{I}_{a}$): $m_{i}\sim\mathrm{Bernoulli}(\pi)$
  • Suffix ($i>L_{c}+Ba$): $m_{i}=1$
6:Sample diffusion step $t$ with probability $\propto\omega_{t}$
7:Apply noise: $\mathbf{z}_{t}\sim q_{t}(\cdot\mid\mathbf{x})$
@@
8:Compute loss only on active block: $\mathcal{L}=-\sum_{i\in\mathcal{I}_{a}}\log p_{\theta}(x_{i}\mid\mathbf{z}_{t},t)$
@@
9:Backpropagate through active block only; treat prefix/suffix as constants
10:Update $\theta$ using gradient $\nabla_{\theta}\mathcal{L}$ with standard optimizer
Implementation details. Blockwise SFT is a drop-in replacement for classical SFT: the only change is a structured mask that preserves the prefix and hides all future tokens. We keep the standard diffusion training recipe (Appendix A.1); each step uses the stochastic gradient estimator in Theorem 3.3, with uniform block sampling $\rho(a)=1/M$ and importance-weighted diffusion steps to optimize—without bias—the variational bound in Theorem 3.2. In practice we use sequence length L=128L{=}128 to balance speed and context; uniform block sampling works well, though difficulty-aware reweighting could further improve efficiency. The method requires no architectural changes and integrates cleanly with existing diffusion LM codebases. By maintaining clean prefixes, hiding suffixes, and focusing loss on the active block, Blockwise SFT aligns training with blockwise inference and removes the systematic bias of classical SFT at comparable compute.
4 Experiments
4.1 Experiment Setup
We train on meta-math/MetaMathQA (Yu et al., 2024) and evaluate on the test splits of openai/gsm8k (Cobbe et al., 2021) and HuggingFaceH4/MATH-500 (Hendrycks et al., 2021), reporting Pass@1 via exact string match. These mathematical reasoning sets align with our blockwise paradigm: each step depends causally on preceding derivations without access to future tokens, mirroring clean-prefix, hidden-future supervision.
Configurations. We compare (1) base (no fine-tuning), (2) Classical SFT with random masking over the full response, and (3) Blockwise SFT with active-block masking. All experiments use GSAI-ML/LLaDA-8B-Instruct with LoRA (Hu et al., 2021), following the recipe of Biderman et al. (2024).
Evaluation Protocols. Recall that Classical SFT aggregates loss over the entire response each step, whereas Blockwise SFT aggregates loss only on the active block (clean prefix, fully hidden suffix). To isolate supervision granularity from raw compute, we report two comparisons: (i) hold model-side compute fixed while changing which tokens receive loss (Equal-FLOPS); and (ii) hold the total number of supervised tokens fixed while allowing the number of optimization steps to differ (Equal-Tokens).
Equal-FLOPS. We fix the FLOP-determining knobs: parameter count, sequence length $L$, batch size, gradient accumulation, and number of optimizer updates. Thus, both methods consume identical forward and backward FLOPs; only the loss-bearing positions differ. For example, with $L_{c}{=}32$, $L_{r}{=}96$, and block size B=32B{=}32, both run the same updates; Classical applies loss over masked positions among $96$ response tokens per step, while Blockwise applies loss only within the sampled active block of $32$ tokens.
Equal-Tokens. We equalize the total number of supervised tokens over the run and index progress by $\tau$ (traversals): for Classical SFT, $\tau$ equals epochs; for Blockwise, $\tau{=}1$ means every sample’s blocks are each supervised once (requiring $\approx L_{r}/B$ epochs). In the same example, Classical supervises $96$ tokens/step while Blockwise supervises $32$; accordingly, Blockwise runs $3{\times}$ more epochs to reach the same $\tau$.
Implementation Details. Bernoulli masking rate $\pi\!\sim\!\mathrm{Uniform}(10^{-3},1)$ is sampled per sequence (Classical) or per active block (Blockwise); if nothing is masked, we mask the last token. Training uses AdamW (Loshchilov & Hutter, 2019) with linear warmup and cosine decay (Smith, 2017) in bfloat16 (Micikevicius et al., 2018), runs on a single A100 (80GB) at $\sim$5.1k tokens/s, and peaks at $\approx$23.5GB memory (deployable on RTX 3090) when the batch size is set to 1. Full hyperparameters are in Appendix A.1.
@@ 4.2 Main Results
Table 1 presents our primary results comparing Blockwise SFT against the original model and five baseline methods under matched compute. Blockwise SFT achieves the highest Pass@1 accuracy on both benchmarks, with a substantial +5.2 point gain over the strongest baseline (Two-Step Loss & Scheduling) on GSM8K and +1.6 points on MATH. Notably, Classical SFT underperforms the base model on MATH (29.6 vs. 31.7), suggesting that naive supervision misalignment can be detrimental. Recent variants offer only marginal improvements, while Blockwise SFT’s aligned supervision delivers consistent and significant gains.
@@
Training Dynamics under Equal-FLOPS and Equal-Tokens. Figs. 3a–b show that Blockwise SFT converges to a lower loss (0.24) than Classical SFT (0.36), indicating more effective learning under aligned supervision. Figs. 4a–b track test Pass@1 throughout training: on GSM8K, Blockwise SFT maintains a $\sim$10-point margin; on MATH, Classical SFT degrades below the base model while Blockwise SFT improves steadily. These dynamics highlight that training–inference alignment is crucial to preserve and enhance reasoning ability.
(a) Classical SFT
(b) Blockwise SFT
Figure 3: Training loss on MetaMathQA under Equal-FLOPS.
(a) GSM8K
(b) MATH
Figure 4: Test Pass@1 during training under Equal-FLOPS.
Fig. 5 evaluates accuracy under a fixed supervised-token budget. Blockwise SFT delivers large gains at $\tau{=}1$ and remains stable through $\tau{=}3$. By contrast, Classical SFT behaves differently across datasets: on GSM8K it shows a small early uptick at $\tau{=}1$ before degrading as $\tau$ increases, whereas on MATH it declines monotonically across traversals. The late-epoch drops further indicate clear overfitting for Classical SFT when the same tokens are revisited over multiple epochs, while Blockwise SFT remains robust under the same budget. Thus, Blockwise SFT’s advantage stems from supervision quality, not quantity.
(a) GSM8K
(b) MATH
Figure 5: Test Pass@1 during training under Equal-Tokens.
@@ 4.3 Block Size Consistency Study
Figure 6 investigates robustness to block size misalignment between training and inference. For each dataset, we train four checkpoints with $B_{\text{train}}\!\in\!\{8,16,32,64\}$ under the Equal-FLOPS protocol (identical sequence length, batch size, and number of optimizer updates). Each checkpoint is then evaluated with semi-autoregressive decoding at $B_{\text{infer}}\!\in\!\{8,16,32,64\}$, producing a $4{\times}4$ grid per dataset. Diagonal cells capture matched granularity; off-diagonal cells impose a controlled mismatch at inference.
In both experimental groups, performance peaks along the diagonal ($B_{\text{train}}{=}B_{\text{infer}}$) and degrades progressively as the mismatch grows. This diagonal dominance indicates that gains stem from aligning supervision granularity with the decoding procedure, rather than from any single block size. Near-diagonal cells degrade gracefully, suggesting robustness to minor misalignment. The factorial design removes compute confounds, attributing the observed pattern to training–inference alignment.
(a) GSM8K
(b) MATH
Figure 6: Test Pass@1 across $B_{\text{train}}$ and $B_{\text{infer}}$ combinations on two datasets.
4.4 Ablation Studies
Prefix Noise Analysis. Figs. 7a–b examine the effect of corrupting prefix blocks during training. Performance degrades progressively as $\pi_{\text{prefix}}$ increases, with the steepest drop at $\pi_{\text{prefix}}=1.0$. This degradation pattern has an important nuance: when block sizes are large, noisy prefixes often have no effect because the first block (which has no prefix) is selected more frequently. This explains why Blockwise SFT maintains reasonable performance even with substantial prefix corruption, because the fast convergence compensates for the effectively reduced training data. The key insight is that maintaining clean prefix context during training mirrors the deterministic prefix available at inference, reinforcing proper conditioning behavior. Implementation details are provided in Appendix A.7.
(a) Prefix noise on GSM8K
(b) Prefix noise on MATH
Figure 7: Prefix noise ablation under Blockwise SFT: (a) GSM8K and (b) MATH. Performance consistently degrades as $\pi_{\text{prefix}}$ increases, highlighting the importance of clean prefixes.
Suffix Leakage Analysis. Figs. 8a–b reveal that suffix visibility is critically detrimental. As $\pi_{\text{suffix}}$ decreases (more future tokens become visible), performance drops sharply. At $\pi_{\text{suffix}}=0$ (fully visible suffix), Blockwise SFT degenerates to Classical SFT performance. This dramatic degradation occurs because visible suffixes fundamentally corrupt the learning objective: instead of learning causal reasoning from prefix to active block, the model learns to extract answers from future context, which is a capability unavailable at inference. This finding underscores that strict future masking is the primary mechanism behind Blockwise SFT’s success, ensuring the model develops genuine reasoning capabilities rather than spurious correlations with leaked information.
(a) Suffix leakage on GSM8K
(b) Suffix leakage on MATH
Figure 8: Suffix leakage ablation under Blockwise SFT: (a) GSM8K and (b) MATH. Making future tokens visible (smaller $\pi_{\text{suffix}}$) sharply harms performance, underscoring the need for strict future masking.
@@ 5 Conclusion
We introduce Blockwise SFT, a simple recipe that aligns supervision with semi-autoregressive, blockwise decoding in discrete diffusion LMs. By freezing clean prefixes, masking only the active block, and strictly hiding future tokens, it removes the contextual shift and dependency leakage of classical SFT. We ground this method with a variational upper bound on blockwise likelihoods, unbiased block/time-sampled gradients, and an analysis of classical SFT’s mismatch bias (Theorems 3.13.2). Empirically, fine-tuning on MetaMathQA with evaluation on GSM8K and MATH yields consistent gains under matched compute and token budgets, state-of-the-art head-to-head results, and diagnostics that track alignment (block-size consistency in §4.3, prefix/suffix ablations in §4.4). The recipe is easy to adopt: no architectural changes, a drop-in loss, and practical single-step estimators (§3.4). Limitations and next steps include adaptive or uncertainty-aware block sizing, combining with preference or RL-style objectives, and extending the alignment principle to other semi- or non-autoregressive generators. Overall, honoring the decoder’s structural constraints in SFT is a core driver of performance in diffusion-based language models.
@@ A.5 Practical Estimators and Weighting
To reduce cost, we may sample a single diffusion step $t$ with probability $\tilde{\omega}_{t}\propto\omega_{t}$ and multiply the per-step gradient by the normalizer $Z=\sum_{s=1}^{T}\omega_{s}$. Formally, letting
@@
which equals $\nabla_{\theta}\,\mathcal{L}_{\text{BW--SFT}}(\theta)$. Finally, using the standard diffusion–ELBO correspondence—that the weighted per-step reconstruction terms act as a faithful surrogate for the t=0t{=}0 blockwise negative log-likelihood (see Theorem 3.2)—we may identify the inner gradient with that of the t=0t{=}0 blockwise NLL; hence the estimator targets $\nabla_{\theta}\,\mathcal{R}_{\text{block}}(\theta)$ up to an overall positive scalar. ∎